# Preventing Model Collapse Under Overparametrization: Optimal Mixing Ratios for Interpolation Learning and Ridge Regression

### v1

@@ Abstract
Model collapse occurs when generative models degrade after repeatedly training on their own synthetic outputs. We study this effect in overparameterized linear regression in a setting where each iteration mixes fresh real labels with synthetic labels drawn from the model fitted in the previous iteration. We derive precise generalization error formulae for minimum-$\ell_{2}$-norm interpolation and ridge regression under this iterative scheme. Our analysis reveals intriguing properties of the optimal mixing weight that minimizes long-term prediction error and provably prevents model collapse. For instance, in the case of min-$\ell_{2}$-norm interpolation, we establish that the optimal real-data proportion converges to the reciprocal of the golden ratio for fairly general classes of covariate distributions. Previously, this property was known only for ordinary least squares, and additionally in low dimensions. For ridge regression, we further analyze two popular model classes – the random-effects model and the spiked covariance model – demonstrating how spectral geometry governs optimal weighting. In both cases, as well as for isotropic features, we uncover that the optimal mixing ratio should be at least one-half, reflecting the necessity of favoring real-data over synthetic. We validate our theoretical results with extensive simulations.
@@ 1 Introduction
Modern AI models are increasingly trained on their own synthetic outputs. However, this practice can lead to model collapse, where prediction performance degrades progressively with iterative re-training on AI generated synthetic data  Shumailov et al. (2024; 2023). The phenomenon has been empirically observed across a wide array of settings Alemohammad et al. (2024); Bertrand et al. (2024); Bohacek & Farid (2023); Briesch et al. (2023); Hataya et al. (2023); Martínez et al. (2023a; b). Motivated by these observations, recent work has rigorously studied model collapse and developed methods to mitigate it in some cases Shumailov et al. (2024); Dohmatob et al. (2024a; b; 2025); Feng et al. (2025); Dey & Donoho (2024); Gerstgrasser et al. (2024); Kazdan et al. (2024). However previous work remains limited to low-dimensional settings or Gaussian features, creating a significant gap in our understanding. This paper breaks this barrier by addressing two fundamental questions: Can mixing fresh real data with synthetic outputs mitigate model collapse in overparametrized problems? What is the optimal mixing ratio that minimizes prediction error in the long run? We provide rigorous answers for overparametrized linear regression, demonstrating how model collapse can be prevented under overparametrization.
@@
Prior work has rigorously studied model collapse across several problem settings, though with important limitations. For low-dimensional Gaussian distribution estimation, Shumailov et al. (2024) shows how repeated use of synthetic data causes the estimated covariance matrix to collapse to zero almost surely, while the sample mean diverges. Similar results hold for linear regression, as established by Dohmatob et al. (2024a), for Gaussian features. Recent work Dohmatob et al. (2024b; 2025); Feng et al. (2025) attributes collapse to a change in scaling laws, with applications to text generation and Gaussian mixture problems. To mitigate model collapse, Gerstgrasser et al. (2024); Dey & Donoho (2024); Kazdan et al. (2024); He et al. (2025) develop a mixing framework, where models are trained on a mixture of real and synthetic data at each iteration. This approach prevents collapse by ensuring that the test error remains bounded even as the number of iterations increase. Crucially, these papers focus exclusively on low-dimensional problems.
@@
We study two broad classes of estimators—minimum-$\ell_{2}$-norm interpolators (Section 3.1) and ridge regression (Section 3.2 and Section 3.3). Modern machine learning (ML) algorithms frequently exhibit implicit regularization Zhang & Yu (2005); Soudry et al. (2018); Gunasekar et al. (2018a)—with appropriate initialization and step sizes, algorithms converge to predictors that achieve remarkable generalization in overparameterized regimes. Implicit regularization has become a cornerstone for understanding why overparameterized models generalize well Bartlett et al. (2020). Within this framework, min-norm interpolators have emerged as a fundamental class of predictors that commonly arise as implicitly regularized limits of gradient-based algorithms Bartlett et al. (2020); Deng et al. (2022); Gunasekar et al. (2018a; b); Liang & Sur (2022); Montanari et al. (2019); Muthukumar et al. (2020); Soudry et al. (2018); Zhang & Yu (2005); Wang et al. (2022); Zhou et al. (2022). At the same time, ridge regression represents a fundamental learning paradigm that has historically provided valuable insights into complex algorithms, often illuminating phenomena observed in deep networks Hastie et al. (2022); Patil et al. (2024). We study these popular classes of predictors for understanding model collapse under overparametrization.
Specifically, we adopt the fresh data augmentation framework from He et al. (2025): given covariates $\mathbf{X}$, at iteration $t$, we generate a new batch of real responses $\mathbf{y}_{t}$ alongside synthetic responses $\widetilde{\mathbf{y}}_{t}$ produced using the estimator from iteration (t1)(t-1). The estimator at iteration $t$ is formed using a weighted mixture of these real and synthetic responses. Concretely, for the min-$\ell_{2}$-norm interpolator and ridge regression with regularization $\lambda>0$, we define
@@
$\widehat{\boldsymbol{\beta}}_{t,\lambda}=(\mathbf{X}^{\top}\mathbf{X}+n\lambda I)^{-1}\mathbf{X}^{\top}\big(w\,\mathbf{y}_{t}+(1-w)\,\widetilde{\mathbf{y}}_{t,\lambda}\big).$ (1.1)
@@
(i) Quantifying the generalization error. In an overparametrized regime (stated precisely in Section 2), we characterize the generalization error as $t\rightarrow\infty$ for both the min-$\ell_{2}$-norm interpolator $\widehat{\boldsymbol{\beta}}_{t}$ (Theorem 3.1)and the ridge estimator $\widehat{\boldsymbol{\beta}}_{t,\lambda}$ (Theorem 3.4). Our results capture the precise dependence of the limiting risk on key problem parameters: the signal strength, feature covariance matrix, regularization level $\lambda$, and mixing proportion $w$. Our work substantially advances the growing literature on interpolation learning and high-dimensional ridge regression Montanari et al. (2019); Deng et al. (2022); Liang & Sur (2022); Wang et al. (2022); Zhou et al. (2022); Bach (2024); Patil et al. (2024); Mallinar et al. (2024); Song et al. (2024), which has previously not examined the impact of synthetic data on these learning problems.
(ii) Characterizing the optimal mixing ratio. We characterize the optimal weight on real labels that minimizes the long-term prediction error across different settings. For min-$\ell_{2}$-norm interpolation, we establish that the asymptotic risk is uniquely minimized at $w^{\star}=1/\varphi$ (the reciprocal of the golden ratio) for any feature covariance matrix with bounded eigenvalues (see equation (2.4)). This phenomenon was previously proved only for ordinary least squares in low-dimensional linear regression with Gaussian features He et al. (2025). For ridge regression, we prove that the risk is log-convex and admits a unique minimum at $w^{\star}$ under several important scenarios: when the covariance is isotropic (Theorem 3.3), or when the covariance follows a spiked model (Section 3.4.2), or when the signal follows a random-effects model (Theorem 3.2). Across all settings, we show $w^{\star}\geq 1/2$, highlighting the necessity of weighting real-data more heavily than synthetic data. Such rigorous analysis of the mixing ratio provides concrete guidance for mitigating model collapse in overparametrized problems, complementing recent theoretical studies that were limited to low-dimensional regression Gerstgrasser et al. (2024); Dey & Donoho (2024).
@@
The rest of the paper is structured as follows. Section 2 formalizes the problem setup and data-augmentation framework. In Section 3, we state our theoretical results – subsection 3.1 considers min-$\ell_{2}$-norm interpolator and subsections 3.2 and 3.3 consider ridge regression for isotropic and anisotropic covariates, respectively. Subsections 3.4.1 and 3.4.2 provide applications of our main results for random effects model and spike covariance models, respectively. Section 4 provides extensive simulations to complement our theoretical findings. Section 5 presents a discussion and future research directions. The proofs of all theoretical results are in the supplementary material.
@@ 2 Problem Setup
2:Initialize: $\widehat{\boldsymbol{\beta}}_{0,\lambda}\leftarrow(\mathbf{X}^{\top}\mathbf{X}+n\lambda I)^{-1}\mathbf{X}^{\top}\mathbf{y}$.
@@
$\widehat{\boldsymbol{\beta}}_{t,\lambda}\leftarrow(\mathbf{X}^{\top}\mathbf{X}+n\lambda I)^{-1}\mathbf{X}^{\top}\big(w\mathbf{y}_{t}+(1-w)\widetilde{\mathbf{y}}_{t,\lambda}\big).$
@@
To capture an overparametrized regime, we assume that $p>n$ with both diverging at a comparable rate, i.e. $p/n\rightarrow\gamma>1$. This means we work with a sequence of problem instances $\{\mathbf{y}(n),\mathbf{X}(n),\boldsymbol{\beta}(n),\varepsilon(n)\}_{n\geq 1}$, with $\mathbf{X}(n)\in\mathbb{R}^{n\times p(n)},\mathbf{y}(n),\varepsilon(n)\in\mathbb{R}^{n},\boldsymbol{\beta}(n)\in\mathbb{R}^{p(n)}$, satisfying equation (2.1) and further assume that
@@
Below we suppress the dependence on $n$ for conciseness. This regime has seen incredible success in modern ML theory in explaining phenomena observed for deep neural networks and other practical algorithms Hastie et al. (2022); Adlam & Pennington (2020); Montanari et al. (2019); Mei & Montanari (2022); Liang & Sur (2022); Cui et al. (2023); Paquette et al. (2024); Emrullah Ildiz et al. (2025); Lu et al. (2025). The regime has also seen enormous utility in high-dimensional statistics, particularly for the development of new theory and methods in challenging contemporary inference problems Bean et al. (2013); El Karoui (2018); Donoho et al. (2009); Bayati & Montanari (2011); Wang et al. (2017); Sur & Candès (2019); Sur et al. (2019); Fan (2022); Li & Sur (2023); Celentano et al. (2023); Jiang et al. (2025).
In the sequel, we operate under the following assumptions on the covariates and errors that are commonly seen in random matrix theory Bai & Silverstein (2010).
@@
(i) The covariates satisfy $\mathbf{X}=\mathbf{Z}{\bm{\Sigma}}^{1/2}$, where $\mathbf{Z}\in\mathbb{R}^{n\times p}$ are random matrices whose entries $Z_{ij}$ are independent random variables with zero mean and unit variance. We further assume that there exists a constant $\tau>0$ by which the $\upsilon$-th moment of each entry is bounded for some $\upsilon>4$:
$\mathbb{E}\left[\left|Z_{ij}\right|^{\upsilon}\right]\leq\tau^{-1}.$ (2.3)
We will assume throughout that $\bm{\Sigma}$} has bounded eigenvalues $s_{1},\ldots,s_{p}$:
@@
(ii) The noises ${\bm{\varepsilon}}_{t},\widetilde{\bm{\varepsilon}}_{t}$ (defined precisely in Algorithm 1) are assumed to have i.i.d. entries with mean 0, variance $\sigma^{2}$, and bounded moments up to any order. That is, for any $\phi>0$, there exists a constant $C_{\phi}$ such that
$\mathbb{E}\left[\left|\varepsilon_{t,1}\right|^{\phi}\right]\leq C_{\phi},\qquad\mathbb{E}\left[\left|\widetilde{\varepsilon}_{t,1}\right|^{\phi}\right]\leq C_{\phi}.$ (2.5)
@@ 3 Main Results
We begin by introducing two measures that feature crucially in the risk of the estimators $\widehat{\boldsymbol{\beta}}_{t,\lambda}$ and $\widehat{\boldsymbol{\beta}}_{t}$. Let $v_{1},\ldots,v_{p}$ denote the eigenvectors of $\bm{\Sigma}$}, i.e., ${\bm{\Sigma}}=\sum_{k=1}^{p}s_{k}v_{k}v^{\top}_{k}$. Define the probability measures:
$\widehat{H}_{p}(x)=\frac{1}{p}\sum_{k=1}^{p}\bm{1}_{s_{k}\leq x},\qquad\widehat{G}_{p}(x)=\frac{1}{\|\boldsymbol{\beta}\|^{2}_{2}}\sum_{k=1}^{p}\langle v_{k},\boldsymbol{\beta}\rangle^{2}\bm{1}_{s_{k}\leq x}.$ (3.1)
@@
Existence and uniqueness of $m(z)$ is well-known (c.f., (Knowles & Yin, 2017, Lemma 2.2)). Further, $m(z)$ is the companion Stieltjes transform of the free convolution of $H$ and $MP_{\gamma}$, where $MP_{\gamma}$ is the Marchenko-Pastur distribution with parameter $\gamma$ Marčenko & Pastur (1967).

### v2

@@ Abstract
Model collapse occurs when generative models degrade after repeatedly training on their own synthetic outputs. We study this effect in overparameterized linear regression in a setting where each iteration mixes fresh real labels with synthetic labels drawn from the model fitted in the previous iteration. We derive precise generalization error formulae for minimum-$\ell_{2}$-norm interpolation and ridge regression under this iterative scheme. Our analysis reveals intriguing properties of the optimal mixing weight that minimizes long-term prediction error and provably prevents model collapse. For instance, in the case of min-$\ell_{2}$-norm interpolation, we establish that the optimal real-data proportion converges to the reciprocal of the golden ratio for fairly general classes of covariate distributions. Previously, this property was known only for ordinary least squares, and additionally in low dimensions. For ridge regression, we further analyze two popular model classes – the random-effects model and the spiked covariance model – demonstrating how spectral geometry governs optimal weighting. In both cases, as well as for isotropic features, we uncover that the optimal mixing ratio should be at least one-half, reflecting the necessity of favoring real-data over synthetic. We study three additional settings: (i) where real data is fixed and fresh labels are not obtained at each iteration, (ii) where covariates vary across iterations but fresh real labels are available each time, and (iii) where covariates vary with time but only a fraction of them receive fresh real labels at each iteration. Across these diverse settings, we characterize when model collapse is inevitable and when synthetic data improves learning. We validate our theoretical results with extensive simulations.
@@ 1 Introduction
Modern AI models are increasingly trained on their own synthetic outputs. However, this practice can lead to model collapse, where prediction performance degrades progressively with iterative re-training on AI generated synthetic data  Shumailov et al. (2024; 2023). The phenomenon has been empirically observed across a wide array of settings Alemohammad et al. (2024); Bohacek and Farid (2023); Briesch et al. (2023); Hataya et al. (2023); Martínez et al. (2023a; b). For instance, Alemohammad et al. (2024); Shumailov et al. (2024; 2023); Martínez et al. (2023a; b) studied settings where a model is iteratively trained on synthetic data generated from the previously trained model, starting from original real data. They also studied settings where a small fraction of the original real data is used in conjunction with the generated synthetic data. They demonstrated that model collapse occurs in both cases. These concerning observations motivated a surge in recent work that studies model collapse rigorously and attempts to develop method that mitigate it where possible Bertrand et al. (2024); Shumailov et al. (2024); Dohmatob et al. (2024a; b; 2025); Feng et al. (2025); Dey and Donoho (2024); Gerstgrasser et al. (2024); Kazdan et al. (2024). However, previous work remains limited to low-dimensional settings or Gaussian features, creating a significant gap in our understanding. This paper breaks this barrier by addressing two fundamental questions: Can mixing fresh real data with synthetic outputs mitigate model collapse in overparametrized problems? What is the optimal mixing ratio that minimizes prediction error in the long run? We provide rigorous answers for overparametrized linear regression, demonstrating how model collapse can be prevented under overparametrization.
@@
Prior work has rigorously studied model collapse across several problem settings, though with important limitations. For low-dimensional Gaussian distribution estimation, Shumailov et al. (2024) shows how repeated use of synthetic data causes the estimated covariance matrix to collapse to zero almost surely, while the sample mean diverges. Crucially, they consider settings where the original real data is fixed and fresh real data is unavailable at each iteration. Similar results hold for linear regression for Gaussian features  Dohmatob et al. (2024a). Recent work Dohmatob et al. (2024b; 2025); Feng et al. (2025) attributes collapse to a change in scaling laws, with applications to text generation and Gaussian mixture problems. To mitigate model collapse, Bertrand et al. (2024); Gerstgrasser et al. (2024); Dey and Donoho (2024); Kazdan et al. (2024); He et al. (2025) develop a mixing framework, where models are trained on a mixture of real and synthetic data at each iteration. This approach prevents collapse by ensuring that the test error remains bounded even as the number of iterations increase. Crucially, these papers focus exclusively on low-dimensional problems.
Beyond the collapse phenomena themselves, several works highlight practical reasons why earlier real labels may not be fully reusable. For legal reasons or otherwise, previously collected labels may need to be explicitly forgotten or discarded: Li et al. (2021) formalize an online-forgetting mechanism for linear regression motivated by storage limitations and regulatory restrictions. Alternatively, older labels become unreliable due to distributional drift Gneiting et al. (2005). Together, these considerations help explain why training procedures that repeatedly combine fresh real data with synthetic outputs arise naturally, and why understanding their long-run behavior is essential.
Beyond the collapse phenomena themselves, several works highlight practical reasons why earlier real labels may not be fully reusable. For legal reasons or otherwise, previously collected labels may need to be explicitly forgotten or discarded: Li et al. (2021) formalize an online-forgetting mechanism for linear regression motivated by storage limitations and regulatory restrictions. Alternatively, older labels become unreliable due to distributional drift Gneiting et al. (2005). Together, with findings in recursive-training settings Alemohammad et al. (2024) that emphasize the necessity of incorporating fresh real data at every iteration, these considerations help explain why training procedures that repeatedly combine fresh real data with synthetic outputs arise naturally, and why understanding their long-run behavior is essential.
@@
We study two broad classes of estimators — minimum-$\ell_{2}$-norm interpolators (Section 3.1) and ridge regression (Section 3.2 and Section 3.3). Modern machine learning (ML) algorithms frequently exhibit implicit regularization Zhang and Yu (2005); Soudry et al. (2018); Gunasekar et al. (2018b) — with appropriate initialization and step sizes, algorithms converge to predictors that achieve remarkable generalization in overparameterized regimes. Implicit regularization has become a cornerstone for understanding why overparameterized models generalize well Bartlett et al. (2020). Within this framework, min-norm interpolators have emerged as a fundamental class of predictors that commonly arise as implicitly regularized limits of gradient-based algorithms Bartlett et al. (2020); Deng et al. (2022); Gunasekar et al. (2018b; a); Liang and Sur (2022); Montanari et al. (2019); Muthukumar et al. (2020); Soudry et al. (2018); Zhang and Yu (2005); Wang et al. (2022); Zhou et al. (2022). At the same time, ridge regression represents a fundamental learning paradigm that has historically provided valuable insights into complex algorithms, often illuminating phenomena observed in deep networks Hastie et al. (2022); Patil et al. (2024). We study these popular classes of predictors for understanding model collapse under overparametrization.
Our work here is motivated by findings in Alemohammad et al. (2024), which suggest that among interesting possible combinations of real and synthetic data mixing, fresh data augmentation is the only scenario where model collapse can be prevented. We work under this fresh data augmentation framework, which we describe below. This framework also appeared in He et al. (2025), though they focused on low-dimensional problems. Given covariates $\mathbf{X}$, at iteration $t$, we generate a new batch of real responses $\mathbf{y}_{t}$ alongside synthetic responses $\widetilde{\mathbf{y}}_{t}$ produced using the estimator from iteration (t1)(t-1). At iteration $t$, the estimator is formed using a weighted mixture of these real and synthetic responses. Concretely, we define
@@
$\widehat{\boldsymbol{\beta}}_{t,\lambda}=(\mathbf{X}^{\top}\mathbf{X}+n\lambda\bm{I})^{-1}\mathbf{X}^{\top}\big(w\,\mathbf{y}_{t}+(1-w)\,\widetilde{\mathbf{y}}_{t,\lambda}\big).$ (1.1)
@@
(i) Quantifying the generalization error. In an overparametrized regime (stated precisely in Section 2), we characterize the generalization error as $t\rightarrow\infty$ for both the min-$\ell_{2}$-norm interpolator $\widehat{\boldsymbol{\beta}}_{t}$ (Theorem 3.1)and the ridge estimator $\widehat{\boldsymbol{\beta}}_{t,\lambda}$ (Theorem 3.4). Our results capture the precise dependence of the limiting risk on key problem parameters: the signal strength, feature covariance matrix, regularization level $\lambda$, and mixing proportion $w$. Our work substantially advances the growing literature on interpolation learning and high-dimensional ridge regression Montanari et al. (2019); Deng et al. (2022); Liang and Sur (2022); Wang et al. (2022); Zhou et al. (2022); Bach (2024); Patil et al. (2024); Mallinar et al. (2024); Song et al. (2024) (see also data selection strategies Rezaei et al. (2025)), which has previously not examined the impact of synthetic data on these learning problems.
(ii) Characterizing the optimal mixing ratio. We characterize the optimal weight on real labels that minimizes the long-term prediction error across different settings. For min-$\ell_{2}$-norm interpolation, we establish that the asymptotic risk is uniquely minimized at $w^{\star}=1/\varphi$ (the reciprocal of the golden ratio) for any feature covariance matrix with bounded eigenvalues (see equation (2.3)). This phenomenon was previously proved only for maximum likelihood estimators in generalized linear models and nonparametric estimation in He et al. (2025). But the focus here was on low-dimensional settings. For ridge regression, we prove that the risk is log-convex and admits a unique minimum at $w^{\star}$ under several important scenarios: when the covariance is isotropic (Theorem 3.3), or when the covariance follows a spiked model (Section 3.4.2), or when the signal follows a random-effects model (Theorem 3.2). Across all settings, we show $w^{\star}\geq 1/2$, highlighting the necessity of weighting real-data more heavily than synthetic data. Such rigorous analysis of the mixing ratio provides concrete guidance for mitigating model collapse in overparametrized problems, complementing recent theoretical studies that were limited to low-dimensional regression Gerstgrasser et al. (2024); Dey and Donoho (2024).
@@
The rest of the paper is structured as follows. Section 2 formalizes the problem setup and data-augmentation framework. In Section 3, we state our theoretical results: Section 3.1 considers the min-$\ell_{2}$-norm interpolator and Sections 3.2 and 3.3 consider ridge regression for isotropic and anisotropic covariates, respectively. Sections 3.4.1 and 3.4.2 provide applications of our main results for random effects and spiked covariance models, respectively. Section 4 and presents simulations to complement our theoretical findings. We conclude with a discussion and outlining of future directions in Section 6. Proofs are deferred to the supplementary material.
@@ 2 Problem Setup
2:Initialize: $\widehat{\boldsymbol{\beta}}_{0,\lambda}\leftarrow(\mathbf{X}^{\top}\mathbf{X}+n\lambda\bm{I})^{-1}\mathbf{X}^{\top}\mathbf{y}$.
@@
$\widehat{\boldsymbol{\beta}}_{t,\lambda}\leftarrow(\mathbf{X}^{\top}\mathbf{X}+n\lambda\bm{I})^{-1}\mathbf{X}^{\top}\big(w\mathbf{y}_{t}+(1-w)\widetilde{\mathbf{y}}_{t,\lambda}\big).$
@@
To capture an overparametrized regime, we assume that $p>n$ with both diverging at a comparable rate, i.e. $p/n\rightarrow\gamma>1$. This means we work with a sequence of problem instances $\{\mathbf{y}(n),\mathbf{X}(n),\boldsymbol{\beta}(n),\bm{\varepsilon}(n)\}_{n\geq 1}$, with $\mathbf{X}(n)\in\mathbb{R}^{n\times p(n)},\mathbf{y}(n),\bm{\varepsilon}(n)\in\mathbb{R}^{n},\boldsymbol{\beta}(n)\in\mathbb{R}^{p(n)}$, satisfying equation (2.1) and further assume that
@@
Below we suppress the dependence on $n$ for conciseness. This regime has seen incredible success in modern ML theory in explaining phenomena observed for deep neural networks and other practical algorithms Hastie et al. (2022); Adlam and Pennington (2020); Montanari et al. (2019); Mei and Montanari (2022); Liang and Sur (2022); Cui et al. (2023); Paquette et al. (2024); Emrullah Ildiz et al. (2025); Lu et al. (2025). The regime has also seen enormous utility in high-dimensional statistics, particularly for the development of new theory and methods in challenging contemporary inference problems Bean et al. (2013); El Karoui (2018); Donoho et al. (2009); Bayati and Montanari (2011); Wang et al. (2017); Sur and Candès (2019); Sur et al. (2019); Fan (2022); Li and Sur (2023); Celentano et al. (2023); Jiang et al. (2025).
In the sequel, we operate under the following assumptions on the covariates and errors that are commonly seen in random matrix theory Bai and Silverstein (2010).
@@
(i) The covariates satisfy $\mathbf{X}=\mathbf{Z}{\bm{\Sigma}}^{1/2}$, where $\mathbf{Z}\in\mathbb{R}^{n\times p}$ are random matrices whose entries $Z_{ij}$ are independent random variables with zero mean and unit variance. We further assume that there exists a constant $\tau>0$ by which the $\upsilon$-th moment of each entry is bounded for some $\upsilon>4$, that is, $\mathbb{E}\left[\left|Z_{ij}\right|^{\upsilon}\right]\leq\tau^{-1}.$ We will also assume that $\bm{\Sigma}$} has bounded eigenvalues $s_{1},\ldots,s_{p}$:
@@
(ii) The noises ${\bm{\varepsilon}}_{t},\widetilde{\bm{\varepsilon}}_{t}$ (defined precisely in Algorithm 1) are assumed to have i.i.d. entries with mean 0, variance $\sigma^{2}$, and bounded moments up to any order. That is, for any $\phi>0$, there exists a constant $C_{\phi}$ such that $\mathbb{E}\left[\left|\varepsilon_{t,1}\right|^{\phi}\right]\leq C_{\phi},\mathbb{E}\left[\left|\widetilde{\varepsilon}_{t,1}\right|^{\phi}\right]\leq C_{\phi}.$
@@ 3 Main Results
We begin by introducing two measures that feature crucially in the risk of the estimators $\widehat{\boldsymbol{\beta}}_{t,\lambda}$ and $\widehat{\boldsymbol{\beta}}_{t}$. Let $\bm{v}_{1},\ldots,\bm{v}_{p}$ denote the eigenvectors of $\bm{\Sigma}$}: ${\bm{\Sigma}}=\sum_{k=1}^{p}s_{k}\bm{v}_{k}\bm{v}^{\top}_{k}$. Define the probability measures:
$\widehat{H}_{p}(x)=\frac{1}{p}\sum_{k=1}^{p}\bm{1}_{s_{k}\leq x},\qquad\widehat{G}_{p}(x)=\frac{1}{\|\boldsymbol{\beta}\|^{2}_{2}}\sum_{k=1}^{p}\langle\bm{v}_{k},\boldsymbol{\beta}\rangle^{2}\bm{1}_{s_{k}\leq x}.$ (3.1)
@@
Existence and uniqueness of $m(z)$ is well-known (c.f., (Knowles and Yin, 2017, Lemma 2.2)). Further, $m(z)$ is the companion Stieltjes transform of the free convolution of $H$ and $MP_{\gamma}$, where $MP_{\gamma}$ is the Marchenko-Pastur distribution with parameter $\gamma$ Marčenko and Pastur (1967).